A positively-charged object with a mass of 0.128 kg oscillates at the end of a spring, generating ELF (extremely low frequency) radio waves that have a wavelength of 6.61 × 107 m. The frequency of these radio waves is the same as the frequency at which the object oscillates. What is the spring constant of the spring?
Answer #1

Taking the speed of light as 3 x 10^8 m/s gives the frequency f as (3*10^8)/(6.61*10^7)

f= 4.538 Hz

So √(k/M) = 2πf

= 2*π*4.538= 28.51

giving k = (28.51*√M)²

k = (28.51*√0.128)^2

k= 104.09 N/m

where M is the mass and k the spring constant.

I hope it was helpful.. please rate me positive..